# Entropic uncertainty under indefinite causal order and input-output direction

### v1

@@ I Introduction
Within the framework of information theory, entropy provides a fundamental and operationally relevant quantifier of uncertainty rather than variance. Whereas variance measures the spread of possible measurement outcomes, entropy characterizes the intrinsic unpredictability of those outcomes. This perspective led to the formulation of entropic uncertainty relations (EURs) [6, 7, 8, 9], in which the total uncertainty associated with two incompatible observables is expressed in terms of the Shannon entropies of their respective measurement outcome distributions. The most well-known EUR was first proposed by Deutsch [10], and following a conjecture by Kraus [11], was later refined by Maassen and Uffink [12], yielding
@@
where $S(\varrho)=-\operatorname{Tr}[\varrho\log_{2}\varrho]$ is the von Neumann entropy, and $S(A|B)=S(\varrho_{AB})-S(\varrho_{B})$ represents the conditional entropy. On the other hand, $S(O|B)=S(\varrho_{OB})-S(\varrho_{B})$, with $O\in\{Q,R\}$, denotes the conditional entropies of the post-measurement states $\varrho_{OB}=\sum_{k}(\Pi_{k}^{O}\otimes\mathbb{I})\varrho_{AB}(\Pi_{k}^{O}\otimes\mathbb{I})$, after a measurement of the observable $O$ is performed on the system $A$, where $\mathbb{I}$ is the identity operator acting on memory $B$. Notice that the lower bound Eq. (2) contains both state independent and state dependent terms. MA-EUR in Eq. (2) was experimentally verified [14, 15]. Furthermore, it was shown to be relevant for numerous subjects in quantum information science, such as quantum key distribution [13, 16], witnessing [17, 18] and sharing entanglement [19], quantum batteries [20], quantum teleportation [21, 22], and more [23, 24].
@@
Since realistic quantum systems are prone to the detrimental effects of the environment [25], MA-EUR has also been investigated under noisy quantum evolutions, which are typically modeled by quantum channels [26]. Given the relevance of MA-EUR in quantum information tasks, understanding the impact of environmental noise on both the total entropic uncertainty and its lower bound constitutes a problem of practical interest. In order to study this problem, different dynamical regimes such as Markovian and non-Markovian, as well as noise settings such as local and global channels, have been considered in the recent literature [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In general, it has been observed that environmental noise tends to increase the total entropic uncertainty by diminishing quantum correlations between the measured system and the quantum memory. However, quantum channels describing the dynamics induced by certain structured or non-Markovian environments can help reduce the uncertainty.
@@
In classical physics, causal relations between events are well defined. Present events are determined by those in the past and, in turn, they influence those in the future. Still, quantum theory admits a more general structure, in which quantum operations may occur without a fixed causal order. It was argued that there exist causally non-separable processes that cannot be described by any definite ordering of events [38, 39, 40, 41, 42, 43]. A paradigmatic example of such a process is the quantum switch [38], which coherently controls the order in which two quantum operations are applied. In essence, quantum switch places two alternative causal orders into superposition, leading to scenarios where the sequence of quantum channels itself becomes indefinite. Quantum switch was experimentally implemented in various platforms [44, 45, 46, 47, 48, 49, 50, 51]. Moreover, the exploitation of indefinite causal order was shown to offer advantages in several areas, such as quantum metrology [52, 53], channel discrimination [54], communication complexity [55], query complexity [56, 57, 58], noisy transmission of information [59, 60, 61, 62, 63, 64, 65, 66, 67], quantum thermodynamics [68, 69, 70, 71], and non-Markovian processes [72, 73, 74].
@@ IV Main Results
In this section, we present our main findings regarding the behavior of MA-EUR in Eq. (2) when the memory qubit $B$ is undergoing noisy dynamics. Before introducing the noise on the memory, let us first elaborate on the operational meaning of the MA-EUR as an uncertainty game between Alice and Bob, who respectively possess the particle $A$ and the quantum memory $B$, thereby providing an interpretation of the entropic bound. In this game, Alice randomly chooses to measure either $Q$ or $R$ on her particle $A$, which is entangled with Bob’s quantum memory $B$. Alice then announces which observable was measured but not the outcome of the measurement. Bob next performs a measurement on his memory qubit $B$ to infer Alice’s measurement result. Bob’s uncertainty (given access to quantum memory $B$) is quantified by the conditional entropies $S(Q|B)$ and $S(R|B)$, whose sum is constrained by the lower bound in Eq. (2). In absence of quantum correlations between $A$ and $B$, the conditional entropy $S(A|B)$ becomes non-negative, which renders the lower bound weaker and provides no advantage over the bound in Eq. (1). Interestingly, for maximally entangled states and complementary observables, both the conditional entropies and the bound vanish, so Bob can predict outcomes of Alice’s measurements with certainty, which happens thanks to the fact that the conditional entropy $S(A|B)$ can take negative values for entangled states [94].
@@ IV.3 Time-flipped Pauli Channels
provides a necessary and sufficient condition for a net decrease in total entropic uncertainty for time-flipped Pauli channels in comparison with the single-use case. We note that, as opposed to the self-switch case, here we are able to obtain an algebraic necessary and sufficient condition thanks to the fact that z-measurement uncertainty is the same for both single use the time-flipped Pauli channels.
@@
Lastly, we consider the complete family of Pauli channels, characterized by $\alpha_{x}$, $\alpha_{y}$ and $\alpha_{z}$, to display the region where the total entropic uncertainty is diminished through the implementation of time-flip superchannel in comparison with the single-use case. The panels in Fig. 5 display the regions where $\Delta U^{\prime}=U_{\text{tf}}-U_{\text{su}}<0$ for four different error probabilities. In each panel, these regions are analytically determined by the condition in Eq. (63). Hence, there is a wide class of Pauli channels, for which the time-flip provides an advantage in reducing the total uncertainty. In particular, the greatest advantage is attained, provided that $\alpha_{z}=0$ and $\alpha_{y}=\min(1,1/2p)$, in which case $\Delta U^{\prime}=-1$ when $\alpha_{y}=1/2p\leq 1$.
V Conclusion
@@
These results were obtained assuming that the complementary measurements on the control qubit are performed in x and z bases. For concreteness, we have also supposed that the bipartite system is initially in one of the Bell states. In fact, our investigation can be straightforwardly extended to the family of Bell-diagonal states since the operations we consider here preserve the Bell-diagonal form of the bipartite state. Therefore, the post-measurement states retain the same structure, and the total uncertainty keep the same functional form as in the Bell state case, up to some simple rescaling factors. Although the choice of a different measurement pair on the control qubit can naturally modify the regions of advantage, the overall treatment remains the same. All in all, our results are not tied to a special choice of initial states or observable pairs. An interesting open direction is to investigate whether similar conclusions hold for non-unital channels, where preservation of Bell-diagonality need not occur and qualitatively new behavior may emerge.

### v2

# arXiv:2511.14192v2[quant-ph] 01 Mar 2026

@@ I Introduction
Within the framework of information theory, entropy provides a fundamental and operationally relevant quantifier of uncertainty rather than variance. Whereas variance measures the spread of possible measurement outcomes, entropy characterizes the intrinsic unpredictability of those outcomes. This perspective led to the formulation of entropic uncertainty relations (EURs) [45, 12, białynicki-birula75, 96], in which the total uncertainty associated with two incompatible observables is expressed in terms of the Shannon entropies of their respective measurement outcome distributions. The most well-known EUR was first proposed by Deutsch [28], and following a conjecture by Kraus [56], was later refined by Maassen and Uffink [63], yielding
@@
where $S(\varrho)=-\operatorname{Tr}[\varrho\log_{2}\varrho]$ is the von Neumann entropy, and $S(A|B)=S(\varrho_{AB})-S(\varrho_{B})$ represents the conditional entropy. On the other hand, $S(O|B)=S(\varrho_{OB})-S(\varrho_{B})$, with $O\in\{Q,R\}$, denotes the conditional entropies of the post-measurement states $\varrho_{OB}=\sum_{k}(\Pi_{k}^{O}\otimes\mathbb{I})\varrho_{AB}(\Pi_{k}^{O}\otimes\mathbb{I})$, after a measurement of the observable $O$ is performed on the system $A$, where $\mathbb{I}$ is the identity operator acting on memory $B$. Notice that the lower bound Eq. (2) contains both state independent and state dependent terms. MA-EUR in Eq. (2) was experimentally verified [57, 75]. Furthermore, it was shown to be relevant for numerous subjects in quantum information science, such as quantum key distribution [13, 74], witnessing [99, 59] and sharing entanglement [48], quantum batteries [87, 88], quantum teleportation [47, 90], and more [26, 93, 65, 97, 94, 95].
@@
Since realistic quantum systems are prone to the detrimental effects of the environment [15], MA-EUR has also been investigated under noisy quantum evolutions, which are typically modeled by quantum channels [69]. Given the relevance of MA-EUR in quantum information tasks, understanding the impact of environmental noise on both the total entropic uncertainty and its lower bound constitutes a problem of practical interest. In order to study this problem, different dynamical regimes such as Markovian and non-Markovian, as well as noise settings such as local and global channels, have been considered in the recent literature [98, 50, 51, 43, 41, 53, 66, 100, 20, 11, 6]. In general, it has been observed that environmental noise tends to increase the total entropic uncertainty by diminishing quantum correlations between the measured system and the quantum memory. However, quantum channels describing the dynamics induced by certain structured or non-Markovian environments can help reduce the uncertainty [11, 6, 50].
@@
In classical physics, causal relations between events are well defined. Present events are determined by those in the past and, in turn, they influence those in the future. Still, quantum theory admits a more general structure, in which quantum operations may occur without a fixed causal order. It was argued that there exist causally non-separable processes that cannot be described by any definite ordering of events [22, 72, 16, 8, 73, 10]. A paradigmatic example of such a process is the quantum switch [22], which coherently controls the order in which two quantum operations are applied. In essence, quantum switch places two alternative causal orders into superposition, leading to scenarios where the sequence of quantum channels itself becomes indefinite. Quantum switch was experimentally implemented in various platforms [77, 80, 39, 18, 35, 36, 92, 81]. Moreover, the exploitation of indefinite causal order was shown to offer advantages in several areas, such as quantum metrology [101, 19, 58, 102], channel discrimination [25], communication complexity [37], query complexity [27, 9, 78], noisy transmission of information [30, 76, 34, 17, 14, 21, 24, 82, 68], quantum thermodynamics [31, 38, 85, 60], and non-Markovian processes [49, 67, 7].
@@ II Pauli Channels
Our focus on the Pauli channels in this work is motivated by three main considerations. First, they constitute a canonical noise model in quantum information, encapsulating various fundamental error mechanisms ranging from biased bit-flips and phase-flips to the isotropic depolarizing channel. Second, their defining property of mapping Pauli operators to Pauli operators makes the resulting dynamics analytically tractable and allows closed-form evaluation of the conditional entropies which appear in the MA-EUR. Finally, the formulation of the quantum time-flip is naturally tailored to bidirectional processes and thus necessitates unitality, a requirement that is automatically satisfied by Pauli channels.
@@ IV Main Results
In this section, we present our main findings regarding the behavior of MA-EUR in Eq. (2) when the memory qubit $B$ is undergoing noisy dynamics. Before introducing the noise on the memory, let us first elaborate on the operational meaning of the MA-EUR as an uncertainty game between Alice and Bob, who respectively possess the particle $A$ and the quantum memory $B$, thereby providing an interpretation of the entropic bound. In this game, Alice randomly chooses to measure either $Q$ or $R$ on her particle $A$, which is entangled with Bob’s quantum memory $B$. Alice then announces which observable was measured but not the outcome of the measurement. Bob next performs a measurement on his memory qubit $B$ to infer Alice’s measurement result. Bob’s uncertainty (given access to quantum memory $B$) is quantified by the conditional entropies $S(Q|B)$ and $S(R|B)$, whose sum is constrained by the lower bound in Eq. (2). In absence of quantum correlations between $A$ and $B$, the conditional entropy $S(A|B)$ becomes non-negative. Thus, despite the lower bound actually still becomes tighter than the bound in relation in Eq. (1), the specific quantum advantage of reducing the total uncertainty enabled by entanglement is lost. [13]. Interestingly, for maximally entangled states and complementary observables, both the conditional entropies and the bound vanish, so Bob can predict outcomes of Alice’s measurements with certainty, which happens thanks to the fact that the conditional entropy $S(A|B)$ can take negative values for entangled states [29].
@@ IV.3 Time-flipped Pauli Channels
provides a necessary and sufficient condition for a net decrease in total entropic uncertainty for time-flipped Pauli channels in comparison with the single-use case. We note that, as opposed to the self-switch case, here we are able to obtain an algebraic necessary and sufficient condition thanks to the fact that z-measurement uncertainty is the same for both single use and time-flipped Pauli channels.
@@
Lastly, we consider the complete family of Pauli channels, characterized by $\alpha_{x}$, $\alpha_{y}$ and $\alpha_{z}$, to display the region where the total entropic uncertainty is diminished through the implementation of time-flip superchannel in comparison with the single-use case. This comparison benchmarks the time-flip against the ordinary single use of the same noise channel, mapping the specific regimes where the implementation of indefinite input-output direction provides a genuine advantage in reducing the total entropic uncertainty. More explicitly, the panels in Fig. 5 display the regions where $\Delta U^{\prime}=U_{\text{tf}}-U_{\text{su}}<0$ for four different error probabilities. In each panel, these regions are analytically determined by the condition in Eq. (63). Hence, there is a wide class of Pauli channels, for which the time-flip provides an advantage in reducing the total uncertainty. In particular, the greatest advantage is attained, provided that $\alpha_{z}=0$ and $\alpha_{y}=\min(1,1/2p)$, in which case $\Delta U^{\prime}=-1$ when $\alpha_{y}=1/2p\leq 1$.
V Conclusion and Discussion
@@
The results presented here can be understood from a common structural property shared by both higher-order quantum processes. In cases of both the quantum switch and the time-flip, the joint output state of the control qubit $A$ and the memory $B$ has a block form in the control basis, where the off-diagonal blocks encode the coherent combination of the two underlying alternatives. When the qubit $A$ is measured in a basis that preserves coherence, e.g., $\sigma_{x}$, these off-diagonal contributions affect the conditional states of qubit $B$ and thus the conditional entropies $S(Q|B)$ and $S(R|B)$. However, measuring $A$ in an incoherent basis, e.g., $\sigma_{z}$, suppresses these contributions. The conceptual difference lies in what alternatives are combined, that is, whereas the quantum switch combines the two causal orders of the channels, the quantum time-flip combines forward versus input-output inverted realizations of the same dynamics. This is basically why both can yield uncertainty reduction while producing distinct advantage landscapes.
Before we conclude, a few last remarks are in order. To start with, it is important to discuss the dependence of our results on the choice of the measurements. In our treatment, we focused on the observables $\sigma_{x}$ and $\sigma_{z}$ for two main reasons. First, they represent the two limiting cases of the interference mechanism. On the one hand, $\sigma_{z}$ is diagonal in the control basis and thus it is effectively blind to the off-diagonal interference terms contributing to the dynamics. On the other hand, $\sigma_{x}$ is purely off-diagonal in the control basis, which lets us fully access the interference terms. Second, this choice renders the dynamics analytically tractable, allowing us to derive exact closed-form conditions for the advantage regions, which would be difficult to obtain for arbitrary observable pairs. While using other incompatible measurement pairs would generally change the quantitative boundaries and thresholds of the advantage regions, the physical mechanism responsible for the reduction of the total entropic uncertainty remains unchanged. Therefore, our findings can be regarded as a representative case study using the canonical observable pair that establishes the physical mechanism of the advantage rather than demonstrating it for a specific choice of measurement pair.
Furthermore, we comment on the dependence of the advantage on the purity of the initial bipartite state. Although we consider a maximally entangled Bell state in our calculations, the same analysis extends straightforwardly to the broad and physically relevant class of Bell-diagonal mixed states, which can be written as
$\varrho_{AB}=\frac{1}{4}\Bigl(\mathbb{I}\otimes\mathbb{I}+\sum_{j=x,y,z}c_{j}\,\sigma_{j}\otimes\sigma_{j}\Bigr),$ (64)
where the coefficients $(c_{x},c_{y},c_{z})$ satisfy the Bell-diagonal positivity constraints [46]. For this family of density operators, the local Pauli channel and the corresponding quantum switch and time-flip constructions built from it preserve the Bell-diagonal structure, and their effect is simply to rescale the coefficients. In the single-use case, the output state coefficients will be scaled as $c_{j}\lambda_{j}$, while for the quantum switch and the quantum time-flip they take the form $c_{j}\kappa_{j}$ and $c_{j}\tau_{j}$, respectively. Consequently, the conditional entropies entering the total entropic uncertainty, $U=S(Q|B)+S(R|B)$, retain the same closed-form expressions as in the Bell state case, with the effective coefficients simply replaced by $c_{j}\lambda_{j}$, $c_{j}\kappa_{j}$, and $c_{j}\tau_{j}$ in the single-use, switch, and time-flip cases, respectively. Thus, the results obtained for the maximally entangled state extend to the broader class of Bell-diagonal states, confirming that the reported advantages reflect a generic feature of the dynamics rather than a consequence of the chosen pure initial state, even though the quantitative advantage landscape is naturally determined by the degree of initial mixedness. An interesting research direction is to investigate whether similar conclusions hold for more general noise models beyond Pauli channels (including unital but non–Pauli-diagonal channels or non-unital dissipative channels), where Bell-diagonal structure need not remain preserved and new behavior may emerge.